OpenStudy (anonymous):

If 4y-x-10=0 and 3x=2y, then xy=

OpenStudy (anonymous):

4y -10 =x put it in 3x=2y 3(4y-10) = 2y 12y-30=2y 10y=30 y=3 3x = 6 x = 2 xy = 2*3 = 6

OpenStudy (anonymous):

This was how I did it. Where did I go wrong? 4y-x=10 3x=2y ; 3x-2y=0 4y-2y-x+3x=10 2y+2x=10 2y+2x/2 =10/2 x=y+5 xy=5